\begin{thebibliography}{}

\bibitem[\protect\citename{Bergsma and Yarowsky}2011]{bergsma11}
Shane Bergsma and David Yarowsky.
\newblock 2011.
\newblock {NADA}: A robust system for non-referential pronoun detection.
\newblock In {\em Anaphora Processing and Applications. Proceedings of the 8th
  Discourse Anaphora and Anaphor Resolution Colloquium, {\em Lisbon, Portugal,
  6-7 October 2011}}, pages 12--23. Springer, Heidelberg.

\bibitem[\protect\citename{Bergsma \bgroup et al.\egroup }2008]{bergsma08}
Shane Bergsma, Dekang Lin, and Randy Goebel.
\newblock 2008.
\newblock Distributional identification of non-referential pronouns.
\newblock In {\em Proceedings of the 46th Annual Meeting of the Association for
  Computational Linguistics: Human Language Technologies, {\em Columbus, Ohio,
  15--20 June 2008}}, pages 10--18.

\bibitem[\protect\citename{Byron and Gegg-Harrison}2004]{byron04}
Donna Byron and Whitney Gegg-Harrison.
\newblock 2004.
\newblock Eliminating non-referring noun phrases from coreference resolution.
\newblock In {\em Proceedings the Discourse Anaphora and Reference Resolution
  Conference}, pages 21--26.

\bibitem[\protect\citename{Clark and Manning}2016a]{clark16a}
Kevin Clark and Christopher~D. Manning.
\newblock 2016a.
\newblock {\GG{20020530}}{I}mproving coreference resolution by learning
  entity-level distributed representations.
\newblock In {\em Proceedings of the 54th Annual Meeting of the Association for
  Computational Linguistics (Volume 1: Long Papers)}, pages 643--653, Berlin,
  Germany, August. Association for Computational Linguistics.

\bibitem[\protect\citename{Clark and Manning}2016b]{clark16b}
Kevin Clark and Christopher~D. Manning.
\newblock 2016b.
\newblock {\GG{20020603}}{D}eep reinforcement learning for mention-ranking
  coreference models.
\newblock In {\em Proceedings of the 2016 Conference on Empirical Methods in
  Natural Language Processing}, pages 2256--2262, Austin, Texas, November.
  Association for Computational Linguistics.

\bibitem[\protect\citename{Denis and Baldridge}2008]{denis08}
Pascal Denis and Jason Baldridge.
\newblock 2008.
\newblock Specialized models and ranking for coreference resolution.
\newblock In {\em Proceedings of the 2008 Conference on Empirical Methods in
  Natural Language Processing, {\em Waikiki, Honolulu, Hawaii, 25--27 October
  2008}}, pages 660--669.

\bibitem[\protect\citename{Evans}2001]{evansrichard01}
Richard Evans.
\newblock 2001.
\newblock Applying machine learning toward an automatic classification of it.
\newblock {\em Literary and Linguistic Computing}, 16(1):45--57.

\bibitem[\protect\citename{Ghaddar and Langlais}2016]{ghaddar16}
Abbas Ghaddar and Philippe Langlais.
\newblock 2016.
\newblock Wiki{C}oref: An {E}nglish coreference-annotated corpus of {W}ikipedia
  articles.
\newblock In {\em Proceedings of the Tenth International Conference on Language
  Resources and Evaluation (LREC 2016)}, Portoro{\v z}, Slovenia, 05/2016.

\bibitem[\protect\citename{Goldberg and Elhadad}2007]{goldberg07svm}
Yoav Goldberg and Michael Elhadad.
\newblock 2007.
\newblock {SVM model tampering and anchored learning: a case study in Hebrew NP
  chunking}.
\newblock {\em {Proceedings of the 45th Annual Meeting of the Association for
  Computational Linguistics}}, pages 224--231.

\bibitem[\protect\citename{Goldberg and Elhadad}2009]{goldberg09}
Yoav Goldberg and Michael Elhadad.
\newblock 2009.
\newblock On the role of lexical features in sequence labeling.
\newblock In {\em Proceedings of the 2009 Conference on Empirical Methods in
  Natural Language Processing, {\em Singapore, 6--7 August 2009}}, pages
  1142--1151. Association for Computational Linguistics.

\bibitem[\protect\citename{Hinton \bgroup et al.\egroup }2012]{hinton12}
Geoffrey~E. Hinton, Nitish Srivastava, Alex Krizhevsky, Ilya Sutskever, and
  Ruslan~R. Salakhutdinov.
\newblock 2012.
\newblock Improving neural networks by preventing co-adaptation of feature
  detectors.
\newblock {\em arXiv preprint arXiv:1207.0580}.

\bibitem[\protect\citename{Hochreiter and Schmidhuber}1997]{hochreiter97}
Sepp Hochreiter and J{\"u}rgen Schmidhuber.
\newblock 1997.
\newblock Long short-term memory.
\newblock {\em Neural Computation}, 9(8):1735--1780.

\bibitem[\protect\citename{Kingma and Ba}2014]{kingma14}
Diederik Kingma and Jimmy Ba.
\newblock 2014.
\newblock Adam: A method for stochastic optimization.
\newblock {\em arXiv preprint arXiv:1412.6980}.

\bibitem[\protect\citename{Lassalle and Denis}2013]{lassalle13}
Emmanuel Lassalle and Pascal Denis.
\newblock 2013.
\newblock Improving pairwise coreference models through feature space hierarchy
  learning.
\newblock In {\em Proceedings of the 51st Annual Meeting of the Association for
  Computational Linguistics (Volume 1: Long Papers), {\em Sofia, Bulgaria, 4--9
  August 2013}}, pages 497--506.

\bibitem[\protect\citename{Lassalle and Denis}2015]{lassalle15}
Emmanuel Lassalle and Pascal Denis.
\newblock 2015.
\newblock Joint anaphoricity detection and coreference resolution with
  constrained latent structures.
\newblock In {\em Proceedings of the 29th Conference on the Advancement of
  Artificial Intelligence, {\em Austin, Texas, 25--30 January 2015}}, pages
  2274--2280.

\bibitem[\protect\citename{Marneffe \bgroup et al.\egroup }2015]{demarneffe15}
Marie-Catherine~de Marneffe, Marta Recasens, and Christopher Potts.
\newblock 2015.
\newblock Modeling the lifespan of discourse entities with application to
  coreference resolution.
\newblock {\em Journal of Artificial Intelligent Research}, 52:445--475.

\bibitem[\protect\citename{Martschat and Strube}2014]{martschat14}
Sebastian Martschat and Michael Strube.
\newblock 2014.
\newblock Recall error analysis for coreference resolution.
\newblock In {\em Proceedings of the 2014 Conference on Empirical Methods in
  Natural Language Processing, {\em Doha, Qatar, 25--29 October 2014}}, pages
  2070--2081.

\bibitem[\protect\citename{Martschat and Strube}2015]{martschat15c}
Sebastian Martschat and Michael Strube.
\newblock 2015.
\newblock Latent structures for coreference resolution.
\newblock {\em Transactions of the Association for Computational Linguistics},
  3:405--418.

\bibitem[\protect\citename{Moosavi and Strube}2016]{moosavi16a}
Nafise~Sadat Moosavi and Michael Strube.
\newblock 2016.
\newblock Search space pruning: A simple solution for better coreference
  resolvers.
\newblock In {\em Proceedings of the 2016 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, {\em San Diego, Cal., 12--17 June 2016}}, pages 1005--1011.

\bibitem[\protect\citename{M{\"u}ller}2006]{mueller06a}
Christoph M{\"u}ller.
\newblock 2006.
\newblock Automatic detection of nonreferential {\em it}\ in spoken multi-party
  dialog.
\newblock In {\em Proceedings of the 11th Conference of the European Chapter of
  the Association for Computational Linguistics, {\em Trento, Italy, 3--7 April
  2006}}.
\newblock 49-56.

\bibitem[\protect\citename{Ng and Cardie}2002]{ng02c}
Vincent Ng and Claire Cardie.
\newblock 2002.
\newblock Identifying anaphoric and non-anaphoric noun phrases to improve
  coreference resolution.
\newblock In {\em Proceedings of the 19th International Conference on
  Computational Linguistics, {\em Taipei, Taiwan, 24 August -- 1 September
  2002}}, pages 730--736.

\bibitem[\protect\citename{Ng}2004]{ng04}
Vincent Ng.
\newblock 2004.
\newblock Learning noun phrase anaphoricity to improve coreference resolution.
\newblock In {\em Proceedings of the 42nd Annual Meeting of the Association for
  Computational Linguistics, {\em Barcelona, Spain, 21--26 July 2004}}, pages
  151--158.

\bibitem[\protect\citename{Ng}2009]{ng09}
Vincent Ng.
\newblock 2009.
\newblock Graph-cut-based anaphoricity determination for coreference
  resolution.
\newblock In {\em Proceedings of Human Language Technologies 2009: The
  Conference of the North American Chapter of the Association for Computational
  Linguistics, {\em Boulder, Col., 31 May -- 5 June 2009}}, pages 575--583.

\bibitem[\protect\citename{Pennington \bgroup et al.\egroup }2014]{glove}
Jeffrey Pennington, Richard Socher, and Christopher~D. Manning.
\newblock 2014.
\newblock Glo{V}e: Global vectors for word representation.
\newblock In {\em Proceedings of the 2014 Conference on Empirical Methods in
  Natural Language Processing, {\em Doha, Qatar, 25--29 October 2014}}, pages
  1532--1543.

\bibitem[\protect\citename{Prince}1992]{prince92}
Ellen~F. Prince.
\newblock 1992.
\newblock The {ZPG} letter: Subjects, definiteness, and information-status.
\newblock In W.C. Mann and S.A. Thompson, editors, {\em Discourse Description.
  Diverse Linguistic Analyses of a Fund-Raising Text}, pages 295--325. John
  Benjamins, Amsterdam.

\bibitem[\protect\citename{Recasens \bgroup et al.\egroup }2013]{recasens13a}
Marta Recasens, Marie-Catherine de~Marneffe, and Christopher Potts.
\newblock 2013.
\newblock The life and death of discourse entities: Identifying singleton
  mentions.
\newblock In {\em Proceedings of the 2013 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, {\em Atlanta, Georgia, 9--14 June 2013}}, pages 627--633.

\bibitem[\protect\citename{Sutskever \bgroup et al.\egroup }2014]{sutskever14}
Ilya Sutskever, Oriol Vinyals, and Quoc~V. Le.
\newblock 2014.
\newblock Sequence to sequence learning with neural networks.
\newblock {\em Advances in Neural Information Processing Systems}, pages
  3104--3112.

\bibitem[\protect\citename{Uryupina}2009]{uryupina09}
Olga Uryupina.
\newblock 2009.
\newblock {Detecting Anaphoricity and Antecedenthood for Coreference
  Resolution.}
\newblock {\em Procesamiento del Lenguaje Natural}, 42.

\bibitem[\protect\citename{van Deemter and Kibble}2000]{vandeemter00}
Kees van Deemter and Rodger Kibble.
\newblock 2000.
\newblock On coreferring: Coreference in {MUC} and related annotation schemes.
\newblock {\em Computational Linguistics}, 26(4):629--637.

\bibitem[\protect\citename{Vinyals \bgroup et al.\egroup }2014]{vinyals14}
Oriol Vinyals, Lukasz Kaiser, Terry Koo, Slav Petrov, Ilya Sutskever, and
  Geoffrey Hinton.
\newblock 2014.
\newblock Grammar as a foreign language.
\newblock {\em arXiv preprint arXiv:1412.7449}.

\bibitem[\protect\citename{Wiseman \bgroup et al.\egroup }2015]{wiseman15}
Sam Wiseman, Alexander~M. Rush, Stuart Shieber, and Jason Weston.
\newblock 2015.
\newblock Learning anaphoricity and antecedent ranking features for coreference
  resolution.
\newblock In {\em Proceedings of the 53rd Annual Meeting of the Association for
  Computational Linguistics (Volume 1: Long Papers), {\em Beijing, China,
  26--31 July 2015}}, pages 1416--1426.

\bibitem[\protect\citename{Wiseman \bgroup et al.\egroup }2016]{wiseman16}
Sam Wiseman, Alexander~M. Rush, and Stuart Shieber.
\newblock 2016.
\newblock Learning global features for coreference resolution.
\newblock In {\em Proceedings of the 2016 Conference of the North American
  Chapter of the Association for Computational Linguistics: Human Language
  Technologies, {\em San Diego, Cal., 12--17 June 2016}}.
\newblock To appear.

\bibitem[\protect\citename{Zhou and Kong}2009]{zhouguodong09}
Guodong Zhou and Fang Kong.
\newblock 2009.
\newblock Global learning of noun phrase anaphoricity in coreference resolution
  via label propagation.
\newblock In {\em Proceedings of the 2009 Conference on Empirical Methods in
  Natural Language Processing, {\em Singapore, 6--7 August 2009}}, pages
  978--986.

\end{thebibliography}

\documentclass[11pt]{article}
\usepackage{eacl2017}
\usepackage{times}
\usepackage{url}
\usepackage{latexsym}
\usepackage{amsthm}
\usepackage{multirow}
\usepackage{tikz}
\usetikzlibrary{positioning,calc,fit}
\usepackage{tkz-graph}
\usepackage{caption}
\usepackage{amsmath}
\usepackage{pgfplots}

\newcommand{\GG}[1]{}
\eaclfinalcopy 

\theoremstyle{definition}
\newtheorem{exmp}{Example}[section]

\title{Use Generalized Representations, But Do Not Forget Surface Features}

\author{Nafise Sadat Moosavi \and Michael Strube\\
  Heidelberg Institute for Theoretical Studies gGmbH \\
  Schloss-Wolfsbrunnenweg 35 \\
  69118 Heidelberg, Germany \\
  {\small \url{{nafise.moosavi|michael.strube}@h-its.org}}}

\date{}

\begin{document}
\maketitle
\begin{abstract}
Only a year ago, all state-of-the-art coreference resolvers were using an extensive amount of surface features.
Recently, there was a paradigm shift towards using word embeddings and deep neural networks, where the use of surface features is very limited. 
In this paper, we show that a simple SVM model with surface features outperforms more complex neural models for detecting anaphoric mentions.
Our analysis suggests that using generalized representations and surface features have different strength that should be both taken into account 
for improving coreference resolution.   
\end{abstract}

\section{Introduction}
\label{sect:intro}
Coreference resolution is the task of finding different mentions that refer to the same entity in a given text.
Anaphoricity detection is an important step for coreference resolution.
An anaphoricity detection module discriminates mentions that are coreferent with one of the previous mentions.
If a system recognizes mention $m$ as non-anaphoric,
it does not need to make any coreferent links for the pairs in which $m$ is the anaphor.

The current state-of-the-art coreference resolvers \cite{wiseman16,clark16a,clark16b}, as well as their anaphoricity detection modules, 
use deep neural networks, word embeddings and a small set of features describing surface properties of mentions.
While it is shown that this small set of features has significant impact on the overall performance \cite{clark16a},
their use is very limited in the state-of-the-art systems in comparison to the embedding features.

In this paper, we first introduce a new neural model for anaphoricity detection that considerably outperforms 
the anaphoricity detection of the state-of-the-art coreference resolver, i.e. deep-coref introduced by \newcite{clark16a}.
However, we show that a simple SVM model that is adapted from our coreferent mention detection approach \cite{moosavi16a}, significantly outperforms the more complex neural models.
We show that the SVM model also generalizes better than the neural model on a new domain other than the CoNLL dataset. 

\section{Discriminating Mentions for Coreference Resolution}
\label{sect:mention}
The recognition of various categories of mentions can be beneficial for coreference resolution.
The detection of the following categories is most common in the literature:
(1) non-referential, (2) discourse-old, and (3) coreferent mentions.
One can also discriminate other categories of mentions 
like mentions that are unlikely to be antecedents or discourse-new mentions \cite{uryupina09}.
However, they are not common in comparison to the above categories.
\subsection{Non-Referential Mentions}
\emph{Non-referential} mentions do not refer to an entity.
These mentions only fill a syntactic position. 
For instance, ``it'' in ``it is raining'' is a non-referential mention.
The approaches proposed by \newcite{evansrichard01}, \newcite{mueller06a}, \newcite{bergsma08}, \newcite{bergsma11} 
are examples of detecting non-referential cases of the pronoun \emph{it}. 
\newcite{byron04} present a more general approach for detecting non-referential noun phrases.
\subsection{Discourse-Old Mentions}
\label{ssect:anaphoricity}
Each mention can be assessed from the point of view of the discourse model \cite{prince92}.
According to the discourse model, a mention may be new, old or inferable.
Mentions which introduce a new entity into the discourse are \emph{discourse-new} mentions.
A discourse-new mention may be a singleton or it may be the first mention of a coreference chain.
For instance, The first ``Plato'' in Example~\ref{anaphoric} is a \emph{discourse-new} mention. 
\begin{exmp}
 \label{anaphoric} 
 \emph{Plato} was a philosopher in Classical Greece. \emph{This philosopher} is the founder of the Academy in Athens. 
  \emph{Plato} died at the age of 81.
\end{exmp}
A \emph{discourse-old} mention refers to an entity that is already evoked in the discourse.
Except for first mentions of coreference chains, other coreferent mentions are \emph{discourse-old}.
For instance, ``this philosopher'' and the second ``Plato'' in Example~\ref{anaphoric} are \emph{discourse-old} mentions.

A mention is \emph{inferable} if the hearer can infer the identity of the mention from another entity that has already been evoked in the discourse.
``the windows'' in Example~\ref{inferable} is an \emph{inferable} mention.
\begin{exmp}
 \label{inferable}
 I walked into \emph{the room}. \emph{The windows} were all open.
\end{exmp}

The detection of discourse-old mentions is commonly referred to as  \emph{anaphoricity detection} 
(e.g.\ \newcite{zhouguodong09}, \newcite{ng09}, \newcite{wiseman15}, \newcite{lassalle15}, inter alia)
while the task of anaphoric mention detection, based on its original definition, is of no use for coreference resolution.
Mentions whose interpretations do not depend on previous mentions are called \emph{non-anaphoric} mentions \cite{vandeemter00}.  
For example, both "Plato"s in Example~\ref{anaphoric} are non-anaphoric.


For consistency with the coreference literature, 
we refer to the task of discourse-old mention detection as anaphoricity detection.

Currently, all the state-of-the-art coreference resolvers learn anaphoricity detection jointly with coreference resolution \cite{wiseman15,wiseman16,clark16a}.
The approaches proposed by \newcite{ng02c}, \newcite{ng04}, \newcite{ng09}, \newcite{zhouguodong09}, \newcite{uryupina09} 
are examples of independent anaphoricity detection approaches.  
\subsection{Coreferent Mentions}
\label{ssect:coreferent}
\newcite{demarneffe15} discriminate mentions as coreferent vs.\ non-coreferent. 
Coreferent mentions are those mentions that appear in a coreference chain.
A non-coreferent mention therefore can be a non-referential noun phrase or a referential noun phrase whose entity is only mentioned once (i.e. singleton).
The proposed approaches of \newcite{recasens13a}, \newcite{demarneffe15}, and \newcite{moosavi16a} discriminate mentions for coreference resolution this way.
\section{Anaphoricity Detection Models}
Anaphoricity detection is the most common approach for discriminating mentions for a coreference resolver.
All of the state-of-the-art coreference resolvers use anaphoricity detection.
In this paper, we compare three different anaphoricity detection approaches: two approaches using neural networks and word embeddings, 
and one using an SVM model and surface features.
\newcite{clark16a} introduce the first neural model.
Since \newcite{clark16a} train their anaphoricity model jointly with the coreference model, 
we refer to this model as the joint model.
We introduce a new anaphoricity detection model as the second neural model using a Long-Short Term Memory (LSTM) network \cite{hochreiter97}.
The third approach is adapted from our state-of-the-art coreferent mention detection \cite{moosavi16a}.
\subsection{Joint Model}
As one of the neural models for anaphoricity detection,
we consider the anaphoricity module of deep-coref\footnote{Available at \url{https://github.com/clarkkev/deep-coref}}, 
the state-of-the-art coreference resolution system introduced by \newcite{clark16a}.
This model has three layers for encoding different types of information regarding a mention.
The first layer encodes the word embeddings of the head, first, last, two previous/following words, and the syntactic parent of the mention.
The second layer encodes the averaged word embeddings of the five previous/following words, all words of the mention, sentence words, and document words.
The third layer encodes the following features of a mention: type, length, position and whether it is embedded in another mention.
The outputs of these three layers are combined into one vector and then get passed through a network with two hidden layers.
This anaphoricity model is trained jointly with the deep-coref coreference model.
\subsection{LSTM Model}
\label{sect:lstm}
In this section we propose a new neural model for anaphoricity detection.
Apart from the properties of the mention itself, 
we consider a limited number of surrounding words.
We first generalize the context of a mention
by removing the mention from the context and replacing it with a special placeholder.
In our experiments, we consider the 10 previous and following words of a mention. 
We concatenate the mention tokens and the head token to the generalized word sequence.
We separate the head and mention tokens in the concatenated sequence using two different placeholders.

The word embeddings of the above sequence are encoded using a bidirectional LSTM.
LSTMs show convincing results on generating meaningful representations  
for various NLP tasks (e.g.\ \newcite{sutskever14} and \newcite{vinyals14}).

We also incorporate a set of surface features
that contains (1) mention type (proper, nominal (definite, indefinite), pronouns (\emph{he, I, it, she, they, we, you})),
(2) string match in the text, (3) string match in the previous context,
(4) head match in the text, (5) head match in the previous context, 
(6) contains tokens of another mention, (7) contains tokens of a previous mention,
(8) contained in another mention, (9) contained in a previous mention, and (10) embedded in another mention. 
These features are concatenated with the output of the bidirectional LSTM and get passed
through one more layer that generates the output.

We also experiment with a more complex model including two different LSTMs for 
encoding mentions and their surrounding words.
We consider longer sequences of previous words 
and an attention mechanism for processing the long sequence.
However, the performance did not improve upon the LSTM model while it considerably increased the training time. 
\subsubsection{Implementation Details}
Hyperparameters are tuned on the CoNLL 2012 development set.
We minimize the cross entropy loss using gradient-based optimization and the Adam update rule \cite{kingma14}. 
We use minibatches of size 50.
A dropout \cite{hinton12} with a rate of 0.3 is applied to the output of LSTM. 
We initialize the embeddings with the 300-dimensional Glove embeddings \cite{glove}.
The size of LSTM's hidden layer is set to 128.
The model is trained in only one epoch.

\subsection{SVM Model}
\label{sect:svm}
Our SVM model introduced in \newcite{moosavi16a}, achieves state-of-the-art results for coreferent mention detection.
This model uses the following set of features: 
lemmas and POS tags of all words of a mention, lemmas and POS tags of the two previous/following words, 
mention string, mention length, mention type (proper, nominal, pronoun, list), string match in the text, and head match in the text.
We use a similar SVM model for anaphoricity detection.
In addition to the features we used for coreferent mention detection, we also add the following features for anaphoricity detection:
string match in the previous context, head match in the previous context, mention words are contained in another mention, 
mention words are contained in a previous mention, mention contains words of another mention, mention contains words of a previous mention.
Similar to \newcite{moosavi16a}, we use an anchored SVM \cite{goldberg07svm} with a polynomial kernel of degree two and remove feature-values that occur less than 10 times.
The use of an anchored SVM with pruning helps the model to generalize better on new domains \cite{goldberg09}.

\section{Performance Evaluation}
\label{conll}
We evaluate the anaphoricity models on the CoNLL 2012 dataset.
It is worth noting that all of the examined anaphoricity detectors in this section use the same mention detection module
and results are reported using system detected mentions.
The performance of the mention detection module is of crucial importance for anaphoricity detection.
Therefore, it is important that the compared anaphoricity detectors use the same mention detection.
\begin{table}[!htbp]
\begin{center}
 \small
 \setlength\tabcolsep{5pt}

 \begin{tabular}{@{}l|lll|lll@{}}
\multicolumn{1}{c}{} & \multicolumn{3}{c@{}}{Non-Anaphoric} & \multicolumn{3}{c}{Anaphoric} \\ \hline
\multicolumn{1}{c}{} & R & P & F1 & R & P &  F1 \\ \hline
   \scriptsize joint \small &- & -& -& $81.81$ & $77.18$ & $79.43$ \\
   \scriptsize LSTM \small & $90.71$ & $92.64$ & $91.66$ & $85.00$ & $81.48$ & $83.20$ \\
   \scriptsize LSTM$^*$ \small & $90.51$ & $87.31$ & $88.88$ & $72.64$ & $78.64$ & $75.52$ \\
   
   \scriptsize {SVM} \small & $92.42$ & $92.61$ & $92.51$ & $84.66$ & $84.30$ & $84.48$  \\ \hline
\end{tabular}
\end{center}
\caption{Results on the CoNLL 2012 test set.\label{table:conll}}
\end{table}

The LSTM model that is described in Section~\ref{sect:lstm} is denoted as \emph{LSTM} in Table~\ref{table:conll}.
In order to investigate the effect of the used surface features, 
we also report the results of the LSTM model without using these features (\emph{LSTM$^*$}).

The following observations can be drawn from the results of Table~\ref{table:conll}:
(1) our LSTM model outperforms the joint model while using less features and being trained independently,
(2) the results of the LSTM$^*$ model is considerably lower than those of LSTM, especially for recognizing anaphoric mentions, 
and (3) the simple SVM model outperforms the neural models in detecting both anaphoric and non-anaphoric mentions.
\subsection{Generalization Evaluation}
In order to investigate the generalization on new domains, we evaluate the LSTM and SVM models on the WikiCoref dataset \cite{ghaddar16}.
The WikiCoref dataset is annotated according to the same annotation guideline as that of CoNLL. 
Therefore, it is an appropriate dataset for performing out-of-domain evaluations when CoNLL is used for training.
For the experiments of Table~\ref{table:wikicoref}, all models are trained on the CoNLL 2012 training data and 
tested on the WikiCoref dataset.

The word dictionary that is used for the LSTM model is built based on the CoNLL 2012 training data.
All words that are not included in this dictionary are treated as out of vocabulary words with randomly initialized word embeddings.
We further improve the performance of LSTM on WikiCoref, by adding the words from the WikiCoref dataset into its dictionary.
The LSTM model trained with this extended dictionary is denoted as \emph{LSTM$^\dagger$} in Table~\ref{table:wikicoref}.
\emph{LSTM$^\dagger$} results are still lower than those of the SVM model while SVM does not use any information from the test dataset. 
Pruning rare lexical features from the training data
along the incorporation of part of speech tags, which are far more generalizable than lexical features,
could explain the generalizability of the SVM model on the new domain.
%
\begin{table}[!htbp]
\begin{center}
\small
\setlength\tabcolsep{5pt}
\begin{tabular}{@{}l|lll|lll@{}}
\multicolumn{1}{c}{} & \multicolumn{3}{c@{}}{Non-Anaphoric} & \multicolumn{3}{c}{Anaphoric} \\ \hline
\multicolumn{1}{c}{} & R & P & F1 & R & P &  F1 \\ \hline
   \scriptsize LSTM \small &  $95.53$ & $89.88$ & $92.62$ & $69.50$ & $84.58$ & $76.31$ \\
   \scriptsize  LSTM$^\dagger$ \small & $93.25$ & $92.78$ & $93.01$ & $79.41$ & $80.57$ & $79.99$  \\   
   \scriptsize {SVM} \small & $93.83$ & $93.05$ & $93.43$ & $80.11$ & $82.07$ & $81.08$ \\
   \hline
\end{tabular}
\end{center}
\caption{Results on the WikiCoref dataset.\label{table:wikicoref}}
\end{table}

\section{Analysis Based on Mention Types}
\label{sect:analysis}
We analyze the output of the LSTM and SVM models on the CoNLL 2012 test set 
to see how well they perform for different types of mentions.
As can be seen from Table~\ref{tab:analysys},
there is not much difference between the performance of LSTM and SVM for recognizing anaphoric pronouns.
SVM detects anaphoric proper names better while LSTM is better at recognizing anaphoric common nouns.

We also analyze the output of LSTM$^*$.
As can be seen, the incorporation of surface features does not affect the detection of anaphoric pronouns very much
while it mainly affects the detection of anaphoric proper names by about 24 percent.

%
In order to see whether the same pattern holds for coreference resolution,
we compare the recall and precision errors of the best coreference system that only uses surface features, 
i.e.\ cort \cite{martschat15c} with singleton features \cite{moosavi16a} \footnote{Available at \url{https://github.com/ns-moosavi/cort/tree/singleton_feature}},
and the state-of-the-art deep coreference resolver, i.e.\ deep-coref \cite{clark16a}.
The comparison of the errors for the CoNLL 2012 test set is shown in Table~\ref{table:error}.
We use the error analysis tool of cort introduced by \newcite{martschat14} for the results of Table~\ref{table:error}.
As can be seen from Table~\ref{table:error},
while deep-coref is significantly better than cort for resolving common nouns and specially pronouns,
its result does not go far beyond that of cort when it comes to resolving proper names.

\begin{table}[!htbp]
\begin{center}
 \small
 \setlength\tabcolsep{5pt}
 \begin{tabular}{@{}l|lll|lll@{}}
\multicolumn{1}{c}{} & \multicolumn{6}{c}{Anaphoric} \\ \hline
\multicolumn{1}{c}{} & R & P & F1 & R & P & F1 \\ \hline
 \multicolumn{1}{c}{} &\multicolumn{3}{c}{Proper names}&\multicolumn{3}{c}{Common nouns} \\ \hline
 \scriptsize LSTM \small & $79.49$ & $82.31$ & $80.88$ & $62.96$ & $65.04$ & $63.99$\\ 
 \scriptsize LSTM$^*$ \small& $47.60$ & $70.09$ & $56.69$ & $46.30$ & $57.75$ & $51.40$\\ 
 \scriptsize SVM \small & $83.80$ & $85.71$ & $84.74$& $52.46$ & $71.98$ & $60.69$\\
\hline
    
\multicolumn{1}{c}{} &\multicolumn{3}{c}{Pronouns} &\multicolumn{3}{c}{Other} \\ \hline
   \scriptsize  LSTM \small & $94.67$ & $85.60$ & $89.91$ & $29.11$ & $63.88$ & $40.00$\\ 
   \scriptsize  LSTM$^*$ \small& $92.67$ & $86.01$ & $89.22$ & $10.13$ & $34.78$ & $15.69$\\
   \scriptsize SVM \small & $95.59$ & $86.29$ & $90.71$& $32.91$ & $76.47$ & $46.02$\\

   \hline
   \end{tabular}
\end{center}
\caption{Anaphoricity results for each mention type on the CoNLL 2012 test set.\label{tab:analysys}}
\end{table}
\begin{table}[!htbp]
\begin{center}
 \small
 \begin{tabular}{l|lll}
& Name & Noun & Pronoun \\ \hline
& \multicolumn{3}{c}{\#Recall Errors} \\ \hline
deep-coref & $1110$ & $1499$ & $1537$ \\
cort & $1145$ & $1638$ & $1655$ \\
\hline
& \multicolumn{3}{c}{\#Precision Errors} \\ \hline
deep-coref & $713$ & $672$ & $1162$ \\
cort & $738$ & $747$ & $1736$ \\
   \end{tabular}
\end{center}
\caption{Coreference error analysis.\label{table:error}}
\end{table}

\section{Discussion}
In this paper we analyze the effect of surface features for anaphoricity detection, 
which is a small but an important step for coreference resolution.
Our analysis shows that surface features, as it was known, are important.
Based on our results, the effects of incorporating surface properties and generalized representations are different for different types of mentions.
These results suggest that apart from a unified model, we should consider different models or at least different features for processing 
different types of mentions
and do not put all the burden on a single model to learn the differences.
The works by \newcite{lassalle13} and \newcite{denis08} are examples of models in which distinct models have been used for various types of mentions.
Besides, our analysis  
shows the importance of surface features for proper names.
Word embeddings are very useful for capturing semantic relatedness.
A coreference resolver that uses word embeddings has a great advantage in better resolution of common nouns and pronouns.
However, the use of surface features in current state-of-the-art coreference resolvers is very limited.
Before going towards using more sophisticated knowledge sources, 
there are still easy victories that can be achieved by incorporating more generalizable surface properties, especially for proper names. 

\section*{Acknowledgments} The authors would like to thank Kevin Clark for his help with the deep-coref software
and Mark-Christoph M{\"u}ller for his helpful comments. 
We would also like to thank the four anonymous reviewers for their detailed comments on an earlier draft of the paper. 
This work has been funded by the Klaus Tschira Foundation, Heidelberg,
Germany. The first author has been supported by a Heidelberg Institute for Theoretical Studies
PhD.\ scholarship.

\bibliographystyle{eacl2017}
\bibliography{mybib}
\end{document}